OpenStudy (anonymous):

Write an equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9

OpenStudy (nali):

y-int is (0,9) so use that to complete the equation y-y= m (x-x) Y-9 = -6 (X-0) Y-9= -6x+6 Y= -6x +6+9 Y= -6x+ 15

OpenStudy (anonymous):

line 3: -6 times 0 is 0 not 6

OpenStudy (nali):

sorry about that Y-9= -6x y= -6x+9

OpenStudy (anonymous):

yes also since you know the y-intercept and the slope you can just plug it into the slope intercept form equation. y=mx+b --> m=slope=-6 and b=y-intercept=9 --> y=-6x+9

OpenStudy (anonymous):

don't forget to give me a medal if you think i did good please.